OpenStudy (anonymous):

hi :) can you help me guys to answer this one? thanks :* 'how to divide polynomial using synthetic division in degree 2?'

OpenStudy (kc_kennylau):

\[\frac{x^2+2x-8}{x-2}:\] \[\begin{array}{c|ccc} &1&2&-8\\ 2&\downarrow&2&8\\\hline &1&4&0 \end{array}\] Ignore the 0 at the end, the quotient is x+4. Therefore x^2+2x-8=(x-2)(x+4)

OpenStudy (anonymous):

i mean, the factor of the polynomial is a quadratic

OpenStudy (kc_kennylau):